solve the linear quadratic system algebraically
y= 5-x^2
y= x-15

Answer by ewatrrr(24785)


Hi
solve the linear-quadratic system algebraically
y= 5-x^2 |Note the use of ^ (uppercase 6)
y= x-15
x-15 = 5-x^2
x^2 + x - 20 = 0
factoring
(x+5)(x-4)=0 Note:SUM of the inner product(5x) and the outer product(-4x) = x
(x+5)= 0 x = -5 y = -20 (-5,-20)
(x-4)=0 x = 4 y = -11 (4,-11)